OpenStudy (anonymous):

3. If g(x) = x2 + 3, find g(4). (1 point) 11 19 16 8 4. If h(x) = 2x + 1, find the value of x if h(x) = 11. (1 point) 5 23 6 10

OpenStudy (anonymous):

Substitute the values. 3. g(x) = x^2 + 3, find g(4) So, 4 = 4^2 + 3. Can you simplify from there? 4. If h(x) = 2x + 1, find the value of x if h(x) = 11 So, 11 = 2x + 1 and solve for x.

OpenStudy (anonymous):

I am not quite sure what you mean by simplify, I am kind of slow with algebra. sorry:(

OpenStudy (anonymous):

Its alright, just solve the equation... For #2 you need to find the value of x. 11 = 2x + 1 Subtract 1 from both sides. 11 - 1 = 2x 12 = 2x divide by 2. x = 6 Do you understand the point?

OpenStudy (anonymous):

Okay thank you so much, Yes i kind of understand now...

OpenStudy (anonymous):

g(4)=4^2+3=16+3=19 for the other question, it's as @jazy did it

OpenStudy (anonymous):

Thank you guys <3

OpenStudy (anonymous):

I have one more question though....

OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20. (1 point) 32 12 14 10

OpenStudy (anonymous):

Yea #1 is as @josiahh did it, my bad. (:

OpenStudy (anonymous):

g(x) is 20 So: g(x) = 2(x − 4) 20 = 2(x - 4) Distribute. 20 = 2x - 8 Add 8. 28 = 2x Divide by 2. x = 14

OpenStudy (anonymous):

same step as @jazy for the other one... now you just need to distribute the 2 first g(x)=2x-8

OpenStudy (anonymous):

I totally understand now! thanks (:

OpenStudy (anonymous):

Glad to hear that. (: